# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

### v4

@@ Abstract
We introduce a geometric framework for the study of Bell nonlocality in Hilbert space, aiming to provide intrinsic measures based solely on the properties of quantum states for a given Bell inequality. Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario. We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of the measurements. Within this framework, we define geometric measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as the trace distance, the Hilbert-Schmidt distance, and the relative entropy. We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios. We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible. In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the local state closest to a Werner state is a Werner state, and analogously for Bell-diagonal states. For each quantum state, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios whenever the corresponding set of local states is accessible.
I Introduction
Quantum nonlocality is generally understood as the impossibility of describing experimental correlations using a local hidden variable model Brunner2014 . A state is considered nonlocal if there exists at least one experimental configuration that leads to a violation of a Bell-type inequality, such as the Clauser-Horne-Shimony-Holt (CHSH) inequality. However, nonlocality is not solely a property of the state itself but also depends on the chosen measurement settings, making its geometric characterization particularly challenging. For instance, it is well known that some states do not violate the CHSH inequality, where both Alice and Bob perform two dichotomic measurements (the 2-2-2-2 scenario), but do violate more general Bell inequalities, such as the 3-3-2-2 scenario.
@@ I Introduction
Furthermore, there exist quantum states that do not violate any Bell inequality individually, but when multiple copies of the state are considered along with joint measurements, the resulting system can exhibit a violation of some Bell inequality Popescu1995 ; Gisin1996 ; Masanes2008 . This phenomenon is known as hidden nonlocality, or superactivation of Bell nonlocality. For instance, it has been shown that a state is distillable if and only if, after applying joint local filtering operations on an arbitrarily large number of copies, it leads to a violation of some Bell inequality Masanes2006 . Another example of nonlocality superactivation has been introduced in a recent work Bavaresco2025 , where the authors consider an appropriate (Bell-local) ancilla, a so-called catalyst, and local operations on the combined system.
A natural approach to quantify the amount of nonlocality of a given state, is to study the geometry of the set of local states within Hilbert space while fixing a particular Bell inequality. Specifically, the distance of a given quantum state from this set provides a meaningful geometric measure of its nonlocality for the chosen protocol. Various distance-based measures, such as trace distance, Hilbert-Schmidt distance, and relative entropy, can be employed to quantify the deviation of a state from the local set.
In this work, we explore the geometric and entropic characterizations of quantum correlations beyond entanglement Horodecki2009 , specifically focusing on Bell nonlocality within Hilbert space. Since entanglement and Bell nonlocality coincide for pure states, geometric approaches to entanglement naturally extend to nonlocality in such cases. However, for mixed states, while Bell nonlocality always implies entanglement, the converse does not necessarily hold, as exemplified by Werner states Werner1989 . This establishes a strict hierarchy between these two forms of quantum correlations in finite-dimensional systems.
In the geometric approach such as the one introduced in the present work, the number of measurements for each party, their possible outcomes, and the related Bell inequalities are uniquely determined. Thereafter, for any given state, one possible choice is to maximize the violation of the inequality over all possible measurements performed by the two parties. If no violation of all the different Bell inequalities occurs, the state is considered local. The corresponding geometric measure of Bell nonlocality will then depend solely on the geometry of the set of local states and on the distance of the state from the set. Alternatively, for each possible type of measurement one can proceed to define the corresponding set of local states and the corresponding, measurement-dependent geometric measure of Bell nonlocality.
In principle, the geometric approach described in the present work is very general, as it applies to any form of quantum nonlocality identified by some type of Bell inequality. What one needs to know is, for each instance, the type of measurement considered, and the corresponding set of local states. As already mentioned, this may turn out to be a very daunting task, given that in general the problem of determining the set of local states associated to a given Bell inequality is NP-hard in the number of measurements and, depending on the Hilbert space dimension, in the number of measurement outcomes.
In the following, we will illustrate the general method by restricting to the CHSH inequality with maximum violation of the inequality over all possible measurements. Notably, for two-qubit systems, a complete parameterization of the set of local states with respect to the CHSH inequality is known, enabling precise geometric analysis. Clearly, this is not the most general case, as other forms of Bell inequalities exist. On the other hand, at least in the case of two qubits, it has been shown that the general problem is NP-hard, and therefore, the analysis must necessarily be limited to a relatively small number of measurements. In this context, it is well known that the number of inequalities other that the CHSH one is very limited Collins2004 . For example, in the case of three possible outcomes on each particle, there is only one other type of Bell inequality that arises and that is distinct from the CHSH one.
We will proceed by first introducing the general theoretical framework before focusing on the violation of the CHSH inequality in two-qubit systems. For such systems, the set of local states is fully characterized, allowing us to illustrate our method effectively. We explicitly compute geometric measures of nonlocality for Bell-diagonal states and Werner states, relying on two key results that we establish: the closest local state to a Werner state remains a Werner state, and the closest local Bell-diagonal state remains a Bell-diagonal state.
@@
The work is organized as follows: In section II, we briefly review the basic concepts and structures in the theory of Bell nonlocality. Then, in section III, we introduce the set of free operations as the set of operations that map the set of local states into itself. In section IV we introduce the geometric and entropic measures of nonlocality and discuss their properties. We address the specific case of the CHSH inequality for two qubits in sections V and VI. Finally, in section VII we discuss our results and comment on some possible upcoming research directions.
II Bell nonlocality for dichotomic observables: CHSH inequality
Quantum nonlocality naturally arises in the context of Bell inequalities Bell1964 ; Clauser1969 ; Laskowski2004 . Consider two observers, Alice and Bob, who are spatially separated. Each of them has a set of observables, with Alice having $n_{a}$ and Bob having $n_{b}$. They do not communicate and independently choose which observable to measure, recording their respective outcomes. The correlations between these observables, determined by their choices and the results obtained, are used to construct Bell inequalities.
These inequalities must hold in a classical framework, where we assume at most local hidden variable models and other commonly accepted assumptions, including choice independence, realism, and the absence of experimental loopholes. In bipartite local hidden variable theories, the conditional probability of Alice choosing observable $x$ and obtaining result $a$, and Bob choosing observable $y$ and obtaining result $b$, satisfies the following relation:
$p(a,b|x,y)=\sum_{\lambda}p(\lambda)p(a|x,\lambda)p(b|y,\lambda),$ (1)
where we can express $p(a,b|x,y)$ in quantum mechanical terms as:
$p(a,b|x,y)={\rm Tr}(\rho^{AB}M_{a|x}\otimes M_{b|y}).$ (2)
Here, $M_{a|x}$ represents a positive operator-valued measure (POVM) associated with outcome $a$ for observable $x$, and similarly, $M_{b|y}$ corresponds to outcome $b$ for observable $y$. It follows directly that for separable states, Eq. (2) implies Eq. (1). Consequently, nonlocal states must be entangled Horodecki1995 ; Werner1989 . However, some entangled states may still exhibit locality Werner1989 ; Masanes2006 .
In the case of two dichotomic observables per observer, the Bell inequality takes the CHSH form:
where $E(x,y)$ denotes the statistical expectation value of the measurement outcomes for observables $x$ and $y$ in a given quantum state. For nonlocal states, Eq. (2) does not hold, leading to a violation of the CHSH inequality, Eq. (3).
III Bell nonlocality: the free operations
@@
The set of local states, i.e., those states that do not violate any Bell inequality, will be denoted in the text as $\mathcal{L}$. In the framework of resource theory Chitambar2019 ; Theurer2017 ; Theurer2021 ; Gallego2015 ; Wolfe2020 ; Dakic2012 ; Pirandola2014 ; Streltsov2017 ; Buscemi2020 ; Contreras2019 ; Gour2008 ; Brandao2013 ; Buscemi2023 ; Barrett2005 ; Schmid2020 ; Kuroiwa2020 ; Gonda2023 ; Bartlett2007 ; DeVicente2014 , local states are also referred to as free states, meaning that they do not possess any resource—in this case, nonlocality.
In the case of two-qubit systems and the CHSH inequality (3), the set of local states can be fully characterized Horodecki1995 ; Masanes2008 and will be discussed in detail in Section V.3 below.
@@
The set of operations that map local states into local states, also known as the set of free operations, consists of local operations and shared randomness (LOSR) Chitambar2019 ; Schmid2020 . These operations can be generally defined as follows:
[…]
@@
where $\Phi^{A}_{\lambda}$ and $\Phi^{B}_{\lambda}$ are completely positive and trace-preserving (CPTP) maps. These transformations cannot take a local state outside the set of local states, as they preserve the hidden variable model. Since LOSR operations are also LOCC (while the converse is not true), they inherit many properties established for the extensively studied class of LOCC transformations. There are a few general properties that free operations should satisfy Theurer2021 ; Chitambar2019 :
@@
IV Geometric and Entropic Measures of Bell nonlocality
IV.0.1 Geometric measures
For a given quantum state $\rho$ we take as measure $\mathcal{M}(\rho)$ of Bell nonlocality any contractive distance (with the exception of the Hilbert-Schmidt distance) $D$ between $\rho$ and the closest local state $\rho_{loc}$:
$\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})\,,$ (5)
where $\sigma\equiv\{\rho_{loc}\}$ denotes the set of local states.
Physically meaningful and mathematically sound geometric measures of quantum nonlocality should satisfy the same key properties enjoyed by the bona fide entropic and geometric measures of quantum entanglement and other quantum correlations (e.g. quantum discord) introduced in the course of time Plenio2007 ; Christandl2004 ; Eisert1999 ; Wootters1998 ; Plenio2005 ; Wei2003 ; Illuminati2008 ; Uyanik2010 ; Geomdiscord2016 ; Geomdiscord2017 ; Vedral1997 ; Vedral1998 ; Baumgratz2014 ; Bengtsson2017 . In particular:
@@ IV.0.1 Geometric measures
    $\mathcal{M}(\rho)=\min_{\rho_{loc}\in\sigma}D(\rho,\rho_{loc})=D(\rho,\rho_{loc}^{min})\geq D(\Phi(\rho),\Phi(\rho_{loc}^{min}))\geq\min_{\rho_{loc}\in\sigma}D(\Phi(\rho),\rho_{loc})=\mathcal{M}(\Phi(\rho)).$ (6)
@@
    $a\mathcal{M}(\rho_{1})+(1-a)\mathcal{M}(\rho_{2})$ $=$ $a\min_{\rho_{loc}\in\sigma}D(\rho_{1},\rho_{loc})+(1-a)\min_{\rho_{loc}\in\sigma}D(\rho_{2},\rho_{loc})$ (7)
    $=$ $aD(\rho_{1},\rho_{loc}^{min})+(1-a)D(\rho_{2},\rho_{loc}^{min^{\prime}})$ (8)
    $\geq$ $D(a\rho_{1}+(1-a)\rho_{2},a\rho_{loc}^{min}+(1-a)\rho_{loc}^{min^{\prime}})$ (9)
    $\geq$ $\min_{\rho_{loc}\in\sigma}D(a\rho_{1}+(1-a)\rho_{2},\rho_{loc})$ (10)
    $=$ $\mathcal{M}(a\rho_{1}+(1-a)\rho_{2}).$ (11)
@@
    where the weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$ and the set $\{\Lambda_{i}\}$ of Kraus operators is such that $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathbb{I}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\sigma$ $\forall\rho^{loc}\in\sigma$. In a sense, property 4 might be considered even more important than property 2, since subselection is a process often readily available under controlled experimental conditions. On the other hand, the former is usually significantly harder to verify than the latter, which is certainly satisfied by all well behaved geometric and entropic measures. It should be noted that convexity (property 3) and monotonicity under subselection (property 4), automatically imply monotonicity under LOSR (property 2). The reverse is obviously not true in general. Focusing only on the relative entropy, proof of property 4 is quite straightforward, in analogy with what has been shown to hold for the resource theory of entanglement Vedral1998 and quantum coherence Baumgratz2014 , and we provide it in Section IV.1.
@@
In the following sections, we introduce geometric and entropic measurements and analyze the properties they satisfy.
IV.0.2 Relative entropy

### v5

# Geometric measures of quantum nonlocality: characterization, quantification, and comparison by distances and operations

@@ Abstract
We introduce a geometric framework for studying Bell nonlocality in Hilbert space, where, for a given quantum state, nonlocality is quantified by the distance between the state and the set of local states. This approach applies to any Bell inequality and any measurement scenario. Whenever the local set is characterized, the proposed nonlocality measure can be computed explicitly.
As a general result, we prove that for any scenario in arbitrary dimension the closest local state to a Werner state is itself a Werner state, and analogously, the closest local state to an isotropic state is again isotropic. In the two-qubit case, we further show that the closest local state to a Bell-diagonal state is Bell-diagonal as well. These structural results are independent of the specific Bell inequality considered, thus revealing intrinsic geometric features of these families of states and providing significant simplifications for computing the proposed measures.
For the Clauser–Horne–Shimony–Holt (CHSH) inequality in two-qubit systems and the Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequality for two qudits of arbitrary finite dimension, we derive explicit geometric measures of nonlocality for Bell-diagonal, Werner, and isotropic states using various distance metrics, including the trace, Hellinger, Hilbert–Schmidt distances, and relative entropy. Furthermore, we prove in all generality that for all scenarios in which the local set is not fully characterized, the geometric measures provide rigorous lower bounds on nonlocality.
1 Introduction
Quantum nonlocality is generally understood as the impossibility of describing experimental correlations using a local hidden variable (LHV) model [1, 2]. A multipartite state is considered nonlocal if there exists at least one experimental configuration, involving non-communicating parties that can at most share randomness, which leads to a violation of a Bell-type inequality.
@@ 1 Introduction
It is clear that a state that appears local with respect to one Bell inequality may still violate a different Bell inequality [3, 4, 5]. It can also happen that a local quantum state can violate a Bell inequality after being preprocessed by local filtering operations, revealing a hidden nonlocality [6, 7]. Furthermore, there exist quantum states that do not violate any Bell inequality individually, but when multiple copies of the state are considered along with joint measurements, the resulting system can exhibit a violation of some Bell inequality. This phenomenon is known as superactivation of Bell nonlocality [8, 9].
Determining whether a state is local, that is, whether its correlations admit a local hidden variable (LHV) description, can be challenging as it requires testing the state against an enormous number of possible experimental scenarios. For example, it has been shown that there exist two-qubit Werner states [10] that do not violate the Clauser–Horne–Shimony–Holt (CHSH) inequality [11] but nevertheless exhibit nonlocality when considering more than two measurements per party [12]. Therefore, for a fixed scenario involving $n$ observers, each allowed to perform $m_{i}$ measurements labeled by $x_{i}$ ($i=1,\dots,n$), and each measurement having $\Delta_{i}$ possible outcomes $a_{i}$, we define the set $\mathcal{L}_{\vec{m},\vec{\Delta}}$ of local states, i.e. the set of states that admit a local hidden variable (LHV) description, as follows:
$p(\vec{a}|\vec{x})=\sum_{\lambda}p(\lambda)\prod_{i=1}^{n}p(a_{i}|\lambda,x_{i}),$ (1)
for any choice of $\vec{a}$ and $\vec{x}$, where
$p(\vec{a}|\vec{x})={\rm Tr}\!\left[\rho\bigotimes_{i=1}^{n}M_{a_{i}|x_{i}}\right].$ (2)
Here, $M_{a_{i}|x_{i}}$ denotes the positive operator-valued measure (POVM) element associated with outcome $a_{i}$ for measurement setting $x_{i}$. It follows directly that, for separable states, Eq. (2) implies Eq. (1). Consequently, any nonlocal state must necessarily be entangled. More generally, the set of local states $\mathcal{L}$ may coincide with the intersection of several sets defined for fixed scenarios.
As mentioned previously, nonlocality expresses the incompatibility of probability distributions obtained in quantum mechanics, or more general nonlocal theories, with those allowed by classical theories. This observation leads naturally to a geometric quantification of quantum nonlocality in terms of the trace distance between classical and quantum probability distributions [13]. Such a geometrical approach provides a direct characterization of Bell nonlocality and allows to introduce well-behaved nonlocality monotones, in close analogy with the resource theory of entanglement.
In this work, inspired by resource-theoretical approaches [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], we introduce a natural geometric quantifier of Bell nonlocality based on the distance of a given quantum state to the set of local states. This approach provides a perspective on nonlocality as an intrinsic property of the state and the considered measurement scenario, rather than as a feature associated solely with the resulting probability distribution. In the case where the convex set of local states is defined as the intersection of all possible $\mathcal{L}_{\vec{m},\vec{\Delta}}$, we obtain geometric measures of nonlocality that depend solely on the quantum state itself.
We introduce different distance-based monotones, showing that they satisfy the same key properties enjoyed by the bona fide entropic and geometric measures of quantum entanglement and other quantum correlations, e.g. quantum discord, introduced in the course of time [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. We prove that for Werner states the closest local state is always a Werner state, and analogously, for isotropic states the closest local state is itself isotropic. In the two-qubit case, we also show that the closest local state to a Bell-diagonal state is Bell-diagonal. These results hold independently of the specific Bell inequality under consideration, and therefore capture intrinsic geometric properties of these families of states.
In practice, the set of local states is known only for specific scenarios. In such cases, the resulting geometric measure not only quantifies nonlocality within the given scenario but also provides a lower bound for more general ones.
@@
The work is organized as follows: In Sect. 2, we introduce the set of free operations as the set of operations that map the set of local states into itself. In Sect. 3 we introduce the geometric and entropic measures of nonlocality and discuss their properties. In Sect. 4 we address the specific case of two qubit-systems, focusing in Sect. 5 on the CHSH inequality. In Sect. 6 we discuss isotropic and Werner states in arbitrary dimensions, including both bipartite and multipartite systems. As relevant examples, we compute the exact expressions of the geometric measures of nonlocality relative to the bipartite Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequality for two-qudit isotropic states and we present a preliminary discussion of the multipartite Mermin inequality in arbitrary dimensions. Finally, in section 7 we review our results and comment on some possible upcoming research directions in the geometric theory of quantum nonlocality.
2 Bell nonlocality: the free operations
@@
The set of local states will be denoted in the text as $\mathcal{L}$. In the framework of resource theory, local states are also referred to as free states, meaning that they do not possess any resource – in this case, nonlocality.
@@
The set of operations that map local states into local states, also known as the set of free operations, consists of local operations and shared randomness (LOSR) [14, 27]. These operations can be generally defined as follows:
$\int_{\lambda}d\lambda\Phi^{A}_{\lambda}\otimes\Phi^{B}_{\lambda},$ (3)
@@
where $\Phi^{A}_{\lambda}$ and $\Phi^{B}_{\lambda}$ are completely positive and trace-preserving (CPTP) maps. These transformations cannot take a local state outside the set of local states, as they preserve the hidden variable model. Since LOSR operations are also LOCC (while the converse is not true), they inherit many properties established for the extensively studied class of LOCC transformations. There are a few general properties that free operations should satisfy [14]:
@@
3 Geometric and Entropic Measures of Bell nonlocality
3.1 Geometric measures
For a given quantum state $\rho$ we take as measure $\mathcal{M}(\rho)$ of Bell nonlocality any contractive distance (with the exception of the Hilbert-Schmidt distance) $D$ between $\rho$ and the closest local state $\rho^{loc}$:
$\mathcal{M}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D(\rho,\rho^{loc})\,$, (4)
where $\mathcal{L}\equiv\{\rho^{loc}\}$ denotes the set of local states.
Physically meaningful and mathematically sound geometric measures of a quantum resource should satisfy the following key properties [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]:
@@ 3.1 Geometric measures
    $\mathcal{M}(\rho)$ $=$ $\min_{\rho^{loc}\in\mathcal{L}}D(\rho,\rho^{loc})=D(\rho,\rho^{loc}_{min})$ (5)
    $\geq$ $D(\Phi(\rho),\Phi(\rho^{loc}_{min}))\geq\min_{\rho^{loc}\in\mathcal{L}}D(\Phi(\rho),\rho^{loc})=\mathcal{M}(\Phi(\rho)).$
@@
    $\hskip-56.9055pta\mathcal{M}(\rho_{1})+(1-a)\mathcal{M}(\rho_{2})$ $=$ $a\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{1},\rho^{loc})+(1-a)\min_{\rho^{loc}\in\mathcal{L}}D(\rho_{2},\rho^{loc})$ (6)
    $=$ $aD(\rho_{1},\rho^{loc}_{min})+(1-a)D(\rho_{2},\rho^{{}^{\prime}loc}_{min})$
    $\geq$ $D(a\rho_{1}+(1-a)\rho_{2},a\rho^{loc}_{min}+(1-a)\rho^{{}^{\prime}loc}_{min})$
    $\geq$ $\min_{\rho^{loc}\in\mathcal{L}}D(a\rho_{1}+(1-a)\rho_{2},\rho^{loc})$
    $=$ $\mathcal{M}(a\rho_{1}+(1-a)\rho_{2}).$
@@
    where the weights $p_{i}={\rm Tr}(\Lambda_{i}\rho\Lambda_{i}^{\dagger})$ and the set $\{\Lambda_{i}\}$ of Kraus operators is such that $\sum_{i}\Lambda_{i}^{\dagger}\Lambda_{i}=\mathds{1}$ and $\Lambda_{i}\rho\Lambda_{i}^{\dagger}\in\mathcal{L}$ $\forall\rho^{loc}\in\mathcal{L}$. In a sense, property 4 might be considered even more important than property 2, since subselection is a process often readily available under controlled experimental conditions. On the other hand, the former is usually significantly harder to verify than the latter, which is certainly satisfied by all well behaved geometric and entropic measures. It should be noted that convexity (property 3) and monotonicity under subselection (property 4), automatically imply monotonicity under LOSR (property 2). The reverse is obviously not true in general. Focusing only on the relative entropy, proof of property 4 is quite straightforward, in analogy with what has been shown to hold for the resource theory of entanglement [44] and quantum coherence [45], and we provide it in Section 3.2.
@@
In the following, we introduce geometric and entropic measurements and analyze the properties they satisfy.
3.1.1 Hilbert-Schmidt distance
The Hilbert-Schmidt distance $D_{\mathrm{HS}}(\rho_{1},\rho_{2})$ between two density operators $\rho_{1}$ and $\rho_{2}$ is defined through the Hilbert-Schmidt norm as follows:
$D_{\mathrm{HS}}(\rho_{1},\rho_{2})\,=\,\|\rho_{1}-\rho_{2}\|_{2}\,=\,\sqrt{{\rm Tr}\left[(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})\right]}\,\,.$ (8)
and the corresponding measure is
$\mathcal{M}_{\mathrm{HS}}(\rho)\,=\,\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{HS}}(\rho,\rho^{loc})\,.$ (9)
The Hilbert-Schmidt (HS) distance has no direct operational meaning and, in general, may not monotonic under all CPTP transformations [46, 47]. For instance, it increases under a trivial extension of local subsystems. On the other hand, it provides easily computable bounds on valid contractive distances that are often more difficult to compute.
3.1.2 Hellinger distance
$D_{\mathrm{He}}(\rho_{1},\rho_{2})=\|\sqrt{\rho_{1}}-\sqrt{\rho_{2}}\|_{2}.$ (10)
$D_{\mathrm{He}}(\rho_{1},\rho_{2})=\sqrt{2-2\sum_{i}\sqrt{e^{(1)}_{i}e^{(2)}_{i}}}=\sqrt{\displaystyle\sum_{i=1}^{4}\bigg(\sqrt{e_{i}^{(1)}}-\sqrt{e_{i}^{(2)}}\bigg)^{2}}.$ (11)
The above is the Euclidean norm in the $\sqrt{e_{i}}$-space and coincides with the classical Hellinger distance quantifying the distance between the two classical probability distributions [48]. The corresponding Hellinger measure of Bell nonlocality is thus defined as
$\mathcal{M}_{\mathrm{He}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{He}}(\rho,\rho^{loc})\,.$ (12)
3.1.3 Bures distance
The Bures distance is defined as [46]
$D_{\mathrm{B}}(\rho_{1},\rho_{2})=\sqrt{2\big[1-\sqrt{F(\rho_{1},\rho_{2}})\big]}\,$, (13)
where $F$ is the Uhlmann fidelity given by [49, 50]
$F(\rho_{1},\rho_{2})=\bigg(\mathrm{Tr}\bigg[\sqrt{\sqrt{\rho_{1}}\rho_{2}\sqrt{\rho_{1}}}\bigg]\bigg)^{2}\,.$ (14)
$\mathcal{M}_{\mathrm{B}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{B}}(\rho,\rho^{loc})\,.$ (15)
3.1.4 Trace distance
The trace distance is defined as
$D_{\mathrm{Tr}}(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}=\frac{1}{2}{\rm Tr}\left[\sqrt{(\rho_{1}-\rho_{2})(\rho_{1}-\rho_{2})}\right]\,$, (16)
$D_{\mathrm{Tr}}(\rho_{1},\rho_{2})=\frac{1}{2}{\rm Tr}|\rho_{1}-\rho_{2}|=\frac{1}{2}\sum_{i}|e_{i}^{1}-e^{2}_{i}|\,.$ (17)
The corresponding trace-distance measure of Bell nonlocality reads
$\mathcal{M}_{\mathrm{Tr}}(\rho)=\min_{\rho^{loc}\in\mathcal{L}}D_{\mathrm{Tr}}(\rho,\rho^{loc}).$ (18)
The trace distance defines the probability of error in the Helstrom measurement which is the optimal measurement to distinguish two quantum states. It is jointly convex and nonincreasing under CPTP maps [46, 50]. As a consequence, the corresponding measure of Bell nonlocality Eq. (18) satisfies monotonicity under LOSR and monotonicity under mixing (convexity).
3.2 Relative entropy